Given consumption function investment government purchases and taxes with the following function: c=90+0.75(Y-T) I=120 G=180 T= 80


find


a)the aggregate demand function?


b)The equilibrium level of income?


c)the equilibrium level of consumption?


d)the equilibrium level of aggregate demand? e)draw the graph of consumption and aggregate demand to show the equilibrium level of income and consumption?

1
Expert's answer
2022-03-25T15:08:27-0400

A)

The aggregate macroeconomic function is presented by:

AD=C+I+G


B)

The equilibrium level of income

Y=C+I+G where C=90+0.75(Y-T),I=120,G=180,t=80 then

Y=90+0.75(y-80)+120+180

Y=90+0,75y-60+120+180

Y-0.75y=330

0.25Y=330

Y=1,320


C)

C=90+0.75(Y-T),T=80,Y=1320

C=90+0.75(1320-80)

C=90+0.75(1240)

C=90+930

C=1020


D)

AD=C+I+G

AD=1020+120+180

AD=1320


E)